# BiFold: Bimanual Cloth Folding with Language Guidance

### v1

@@ Abstract
Cloth folding is a complex task due to the inevitable self-occlusions of clothes, their complicated dynamics, and the disparate materials, geometries, and textures that garments can have. In this work, we learn folding actions conditioned on text commands. Translating high-level, abstract instructions into precise robotic actions requires sophisticated language understanding and manipulation capabilities. To do that, we leverage a pre-trained vision-language model and repurpose it to predict manipulation actions. Our model, BiFold, can take context into account and achieves state-of-the-art performance on an existing language-conditioned folding benchmark. Given the lack of annotated bimanual folding data, we devise a procedure to automatically parse actions of a simulated dataset and tag them with aligned text instructions. BiFold attains the best performance on our dataset and can transfer to new instructions, garments, and environments.
@@ I INTRODUCTION
Figure 1: Overview: This paper introduces a language-conditioned model that predicts actions for bimanual cloth folding from an observation of the cloth state and a text instruction. To address the lack of datasets for this task, we present an annotation pipeline that augments an existing dataset of human demonstrations in a simulated environment with aligned language instructions.
@@
  • A new model that leverages foundational vision and language models to learn cloth folding manipulations taking previous actions into account (Section III-A).
  • A procedure to automatically generate language annotations aligned to manipulation actions for existing datasets of folding demonstrations (Section III-B).
@@ II-B Goal-conditioned cloth folding
There are other options to specify folding actions other than images and text. For example, SpeedFolding [6] can take enumerate user-specified folding lines. These lines are drawn on top of an image of the initial configuration of the garment and used along a segmentation mask to define pick and place positions.
@@ III-A Learning language-conditioned pick and place positions
Given an RGB image observation $\mathbf{o}_{t}$ and a language instruction $\ell_{t}$, we want to design a policy $\pi_{\theta}(\mathbf{a}_{t}|\ell_{t},\mathbf{o}_{t})$ parametrized by $\theta$ to obtain manipulation actions $\mathbf{a}_{t}$ at step $t$. In this work, we constrain $\mathbf{a}_{t}$ to a quasi-static pick-an-place manipulation. We include the details of the manipulation primitive in the project page. While other works incorporate other actions such as flings and drags [6], folding a garment can be done with only pick-and-place operations, which are the only ones considered in the unimanual and bimanual datasets used in this work. Moreover, quasi-static actions simplify cloth manipulations since rapid movements are more unpredictable, making it hard to obtain optimal folding results. Instead, the robot can better control the cloth’s shape and prevent unwanted wrinkles by applying slow and controlled movements.
We leverage the SigLIP model [16] to obtain high-level representations of $\mathbf{o}_{t}$ and $\ell_{t}$. SigLIP is a pre-trained contrastive vision-language model similar to CLIP [10] which consists of image and text transformer-based encoders that provide aligned features. Compared to other vision backbones [18, 19, 20, 21] it offers better visual representations [22], presents better 3D awareness [23], offers superior visually-situated text understanding capabilities [24, 25], and suffers a lower sim-to-real gap [26].
We fuse the image and language features by concatenating them and applying attention layers [12]. This is an increasingly popular approach applied to innumerable problems in different domains, including but not limited to tens of modalities [27, 28], robotics [29, 30], and even apparently very different domains such as atmospheric data [31]. By using SigLIP, the features are aligned before the fusion step similarly to ALBEF [32].
We then feed the processed the processed observation tokens as input to convolutional-based decoders. While the decoder can also be based on transformers [21], we found that these yield patch artifacts in low data regimes. The decoders yield probability distributions over the observation space used to sample pick and place locations. Given that the pick position has to fall on the garment, we use a segmentation mask to constrain the support of the predicted distributions. Given the output probabilities, we greedily sample the most probable position to obtain the pick-and-place action $\mathbf{a}_{t}$.
@@
Folding a cloth requires at least two steps. While the text instructions indicate the pick and place positions in natural language, reducing the inherent ambiguity due to the usual garment symmetry, some late manipulation steps may be ill-defined given the current observation. For example, observe how it is impossible to tell top and bottom apart in Fig. 3(b), motivating the use of a policy $\pi_{\theta}(\mathbf{a}_{t}|\ell_{t},\mathbf{o}_{t},\mathbf{o}_{t-1},\dots,\mathbf{o}_{t-H})$ that uses a context of $H$ previous observations. To incorporate context, we encode the previous observations using the same encoders and add a learned positional encoding element-wise for the model to differentiate the time steps. We set a fixed context size $H=3$ and use attention masking to filter out empty tokens if the context does not fill the maximum allowed observation horizon. The pick and place decoders still take the tokens corresponding to $\mathbf{o}_{t}$ as input, and the output tokens of $\mathbf{o}_{t-1}\dots,\mathbf{o}_{1}$ and $\ell_{t}$ are discarded. We depict the architecture of our model in Fig. 2, which we dub BiFold.
@@ III-B Creating aligned language instructions
Preprocessing: All garments in VR-Folding have a green pattern for the exterior of the mesh and an orange color for the interior, hindering generalization to other textures. For this reason, we re-render RGB-D images using the simulation meshes and the original textures of the CLOTH3D dataset [33]. VR-Folding dataset contains some instances in which the simulator became unstable and yielded unrealistic clothes, resulting in sharp and spiked garments. We filter out the divergent meshes by thresholding the ratio between the maximum z-score of the edge length distributions of the simulation mesh and the Normalized Object Canonical Space (NOCS) mesh [34].
@@ IV EXPERIMENTS
TABLE II: Bimanual results in test images.
@@
TABLE III: Bimanual results in SoftGym.
@@ IV-A Experimental setup
We train our models with a batch size of 24 using Adam [36] with a learning rate $10^{-4}$ for 100 epochs. We use LoRA [17] with rank 8, $\alpha=32$ and dropout 0.01. The transformer encoder has 8 blocks and 16 heads. We use the SigLIP [16] model with patch size 16. We perform SE(3) augmentations to the input observations during training and standard image normalization. For our dataset, we use 90% of samples for training and the rest for testing.
Simulation evaluation: We load the cloth meshes to the SoftGym simulator [37] and execute a pick and place primitive with one or two grippers. The unimanual dataset [5] is obtained with SoftGym and can be simulated straightforwardly. To simulate the bimanual dataset, we transform the workspace of the VR data collection setup to the same one used for the unimanual dataset and make sure that the meshes in the initial configuration do not diverge due to unsatisfied constraints.
Real world evaluation: We obtain zenithal view RGB-D images with an Azure Kinect camera. We use SAM [38] for obtaining segmentation masks of the cloth from RGB images. We evaluate our model on different clothes. See more details in the project page. Similarly to previous work [5, 3, 39, 40], we assume that the cloth starts from a flattened state as this is the initial state for all the samples of the training dataset. Note that the dataset contains some refinement actions on states similar to an initial crumpled configuration. However, the language instruction in this case specifies which part to correct and is never the first folding action. If the initial cloth is crumpled, we can use a flattening model [41, 42] to achieve the desired configuration before starting the folding task.
@@ IV-C Bimanual cloth folding
We present the manipulation results on simulation in Table III and a qualitative folding rollout that concatenates several actions in Fig. 5b. We train the models with data from a different renderer, so the sim-to-sim gap affects the performance. Moreover, our dataset contains many more meshes than the unimanual one, and some of them have more than twice the number of vertices of any in the dataset by Deng et al. [5]. For this reason, we report the error, mean Intersection Over Union (IoU), and the success defined by those instances for which the IoU exceeds 80%. Similarly to the image metrics, we can see that BiFold consistently outperforms the baseline. One of the reasons for the drop in performance regarding Table III is the use of human folding demonstrations, which introduces considerably more variability in the demonstrations than using scripted policies on simulation and in some cases are suboptimal (Fig. 5a (i)). Another reason is that Deng et al. [5] obtain pick positions by manually defining some vertices in the simulation mesh, greatly restricting the possible starting positions of the manipulation action. However, human demonstrators can grab any part of the garments. To showcase how BiFold transfers to real-world observations, we perform a qualitative evaluation, which is presented in Fig. 5c. Despite the difference in image appearances, lighting, and unseen garments, BiFold can predict coherent pick-and-place actions.
@@
Figure 5: Qualitative examples: Action predictions obtained with our model. Pick and place actions for the right and left grippers are represented as the origin and endpoints of arrows: red and green for ground truth, and blue and light blue for our model, respectively.
@@ V CONCLUSIONS
BiFold learns folding subactions and we perform complete folds as in Fig. 5b by following a predefined list of ordered instructions, but future work could use a Large Language Model (LLM)-based planner to break down the task into steps [46]. Even if BiFold predicts correct actions, the folding action can fail due to poor simulator physics [47] or incorrect grasps. A natural next step could be to not rely on primitives and directly predict robot actions, e.g., use a diffusion policy [48] and condition it with the predicted pick and place actions [49]. While the NOCS approach is efficacious for CLOTH3D cloth assets [33], it may fail for clothes where topologies are too diverse in a given category (e.g., designer clothes with significant deviations from standard shapes). In this case, we could rely on distilled language features [50].
@@ ACKNOWLEDGMENT
Oriol is supported by project SGR 00514 (Departament de Recerca i Universitats de la Generalitat de Catalunya), and acknowledges travel support from ELISE (GA no 951847).
@@ A-B Language annotations
When annotating bimanual actions using NOCS coordinates (depicted in Fig. 8 for different clothes), it is usual that the left and right pickers use different semantic locations, e.g., the left picker grabs the top right part, and the right picker grabs the bottom right part. In this case, we can infer that the common objective is to hold the right part of the garment, but one cannot trivially resolve many other situations. To do that, we designed a heuristics detailed in Algorithm 1, which outputs a common semantic location considering the positions of the left and right pickers given the context of the action.
@@ C-C Failures
Appendix D Related work
D-A Limitations of previous approaches
@@
Appendix E Limitations and future work
E-A Dataset extension
@@ E-B Modeling and evaluation
Two of the baselines used in this paper [5, 3] use SoftGym as a simulator [37], while the other [4] relies on PyBullet [55]. However, according to the benchmark performed by Blanco-Mulero et al. [47], the cloth physics of these simulators is inaccurate. Among all the simulators tested in the benchmark, MuJoCo [56] is chosen as the one achieving closer dynamics.
Even if the pick and place positions are perfect, the success metric also considers errors from the manipulator. While simulators allow the grasp mesh vertices and hence achieve an ideal grasp, deploying on real robots introduces additional errors due to incorrect grasps or workspace limitations. This can be seen in some examples in simulation, and we include one of them in Fig. 27.

### v2

@@ Abstract
Cloth folding is a complex task due to the inevitable self-occlusions of clothes, their complicated dynamics, and the disparate materials, geometries, and textures that garments can have. In this work, we learn folding actions conditioned on text commands. Translating high-level, abstract instructions into precise robotic actions requires sophisticated language understanding and manipulation capabilities. To do that, we leverage a pre-trained vision-language model and repurpose it to predict manipulation actions. Our model, BiFold, can take context into account and achieves state-of-the-art performance on an existing language-conditioned folding benchmark. To address the lack of annotated bimanual folding data, we introduce a novel dataset with automatically parsed actions and language-aligned instructions, enabling better learning of text-conditioned manipulation. BiFold attains the best performance on our dataset and demonstrates strong generalization to new instructions, garments, and environments.
Figure 1: Overview: This paper introduces a language-conditioned model that predicts actions for bimanual cloth folding. Leveraging an observation of the cloth state and a text instruction, the model generates probability distributions for the pick and place positions of the left and right arms. To address the lack of datasets for this task, we present an annotation pipeline that augments an existing dataset of human demonstrations in a simulated environment with aligned language instructions.
@@ I INTRODUCTION
  • A new model that leverages a foundational vision-language backend to learn cloth folding manipulations taking previous actions into account (Section III-A).
  • A novel dataset of language-aligned bimanual cloth folding actions obtained with an automatic pipeline that could be applied to other datasets (Section III-B).
@@ II-B Goal-conditioned cloth folding
There are other options to specify folding actions other than images and text. For example, SpeedFolding [6] can take enumerated user-specified folding lines drawn on top of an image of the initial configuration of the garment and used along a segmentation mask to define pick and place positions.
@@ III-A Learning language-conditioned pick and place positions
Given an RGB image observation $\mathbf{o}_{t}$ and a language instruction $\ell_{t}$, we want to design a policy $\pi_{\theta}(\mathbf{a}_{t}|\ell_{t},\mathbf{o}_{t})$ parametrized by $\theta$ to obtain manipulation actions $\mathbf{a}_{t}$ at step $t$. In this work, we constrain $\mathbf{a}_{t}$ to a quasi-static pick-an-place manipulation (see the project page for details of the manipulation primitive). While some prior works incorporate additional actions such as flings and drags [6], our approach focuses solely on pick-and-place operations, as these are sufficient for garment folding and are the only actions present in the unimanual and bimanual datasets used to train the policy. Moreover, quasi-static actions simplify cloth manipulations since rapid movements are more unpredictable, making it hard to obtain optimal folding results. Instead, the robot can better control the cloth’s shape and prevent unwanted wrinkles by applying slow and controlled movements.
We leverage the SigLIP model [16] to obtain high-level representations of $\mathbf{o}_{t}$ and $\ell_{t}$. SigLIP is a pre-trained contrastive vision-language model similar to CLIP [10] which consists of image and text transformer-based encoders that provide aligned features. Compared to other vision backbones [18, 19, 20, 21] it offers better visual representations [22], presents better 3D awareness [23], offers superior visually-situated text understanding capabilities [24, 25], and suffers a lower sim-to-real gap in robotics setups [26].
We fuse the image and language features by concatenating them and applying attention layers [12]. This is an increasingly popular approach applied to multimodal models [27, 28] and robotics [29, 30]. By using SigLIP, the features are aligned before the fusion step similarly to ALBEF [31].
We then feed the processed observation tokens as input to convolutional-based decoders. While the decoder can also be based on transformers [21], we found that these yield patch artifacts in low data regimes. The decoders yield probability distributions over the observation space used to sample pick and place locations. Given that the pick position has to fall on the garment, we use a segmentation mask to constrain the support of the predicted distributions. Given the output probabilities, we greedily sample the most probable position to obtain the pick-and-place action $\mathbf{a}_{t}$.
@@
Folding a cloth requires at least two steps. While the text instructions indicate the pick and place positions in natural language, reducing the inherent ambiguity due to the usual garment symmetry, some late manipulation steps may be ill-defined given the current observation. For example, observe how it is impossible to tell top and bottom apart in Fig. 3(b), motivating the use of a policy $\pi_{\theta}(\mathbf{a}_{t}|\ell_{t},\mathbf{o}_{t},\mathbf{o}_{t-1},\dots,\mathbf{o}_{t-H})$ that uses a context of $H$ previous observations. To incorporate context, we encode the previous observations using the same encoders and add a learned positional encoding element-wise for the model to differentiate the time steps. We set a fixed context size $H=3$, as over 95% of the bimanual dataset samples consist of three or fewer actions. This choice balances capturing sufficient historical context while maintaining efficiency and being enough for longer manipulations. We use attention masking to filter out empty tokens if the context does not fill the maximum allowed observation horizon. The pick and place decoders still take the tokens corresponding to $\mathbf{o}_{t}$ as input, and the output tokens of $\mathbf{o}_{t-1}\dots,\mathbf{o}_{1}$ and $\ell_{t}$ are discarded. We depict the architecture of our model in Fig. 2, which we dub BiFold.
@@ III-B Creating aligned language instructions
Preprocessing: All garments in VR-Folding have a green pattern for the exterior of the mesh and an orange color for the interior, hindering generalization to other textures. For this reason, we re-render RGB-D images using the simulation meshes and textures of the CLOTH3D dataset [32]. The VR-Folding dataset contains some instances in which the simulator became unstable and yielded unrealistic clothes, resulting in sharp and spiked meshes. We filter these out by thresholding the ratio between the maximum z-score of the edge length distributions of the simulation mesh and the Normalized Object Canonical Space (NOCS) mesh [33].
@@ IV EXPERIMENTS
TABLE II: Bimanual results on test images.
@@
TABLE III: Bimanual results on SoftGym.
@@ IV-A Experimental setup
We train our models with a batch size of 24 using Adam [35] with a learning rate $10^{-4}$ for 100 epochs. We use LoRA [17] with rank 8, $\alpha=32$ and dropout 0.01. The transformer encoder has 8 blocks and 16 heads. We use the SigLIP [16] model with patch size 16. We perform SE(3) augmentations to the input observations during training and standard image normalization. For our dataset, we use 90% of samples for training and the rest for testing. The model is trained on a single NVIDIA GeForce RTX 3090 GPU.
Simulation evaluation: We load the cloth meshes to the SoftGym simulator [36] and execute a pick and place primitive with one or two grippers. The unimanual dataset [5] is obtained with SoftGym and can be simulated straightforwardly. To simulate the bimanual dataset, we transform the workspace of the VR data collection setup to the one used for the unimanual dataset and make sure that meshes in the initial configuration do not diverge due to unsatisfied constraints.
Real world evaluation: We obtain zenithal view RGB-D images of different clothes with an Azure Kinect camera. We use SAM [37] for obtaining segmentation masks of the cloth from RGB images. See more details in the project page. Similarly to previous work [5, 3, 38, 39], we assume that the cloth starts from a flattened state as this is the initial state for all the samples of the training dataset. Note that the dataset contains some refinement actions on states similar to an initial crumpled configuration. However, the language instruction in this case specifies which part to correct and is never the first folding action. If the initial cloth is crumpled, we can use a flattening model [40, 41] to achieve the desired configuration before starting the folding task.
@@ IV-C Bimanual cloth folding
We present the manipulation results on simulation in Table III and a qualitative folding rollout that concatenates several actions in Fig. 5b. We train the models with data from a different renderer, so the sim-to-sim gap affects the performance. Moreover, our dataset contains many more meshes than the unimanual one, and some of them have more than twice the number of vertices of any in the dataset by Deng et al. [5]. For this reason, we report the error, mean Intersection Over Union (IoU), and the success defined by those instances for which the IoU exceeds 80%. Similarly to the image metrics, we can see that BiFold consistently outperforms the baseline. One of the reasons for the drop in performance regarding Table III is the use of human folding demonstrations, which introduces considerably more variability in the demonstrations than using scripted policies on simulation and in some cases are suboptimal (Fig. 5a (i)). Another reason is that Deng et al. [5] obtain pick positions by manually defining some vertices in the simulation mesh, greatly restricting the possible starting positions of the manipulation action. However, human demonstrators can grab any part of the garments. To showcase how BiFold transfers to real-world observations, we perform an offline qualitative evaluation on test images, which is presented in Fig. 5c. Despite the difference in image appearances, lighting, and unseen garments, BiFold can predict coherent actions.
@@
Figure 5: Qualitative examples: Action predictions obtained with our model. Pick and place actions for right and left are represented as the origin and endpoints of arrows: red and green for ground truth, and blue and light blue for our model.
@@ V CONCLUSIONS
BiFold learns folding subactions and we perform complete folds as in Fig. 5b by following a predefined list of ordered instructions, but future work could use a Large Language Model (LLM)-based planner to break down the task into steps [45]. Even if BiFold predicts correct actions, the folding action can fail due to poor simulator physics [46] or incorrect grasps. A natural next step could be to not rely on primitives and directly predict robot actions, e.g., use a diffusion policy [47] and condition it with the predicted pick and place actions [48]. While the NOCS approach is efficacious for CLOTH3D cloth assets [32], it may fail for clothes where topologies are too diverse in a given category (e.g., designer clothes with significant deviations from standard shapes). In this case, we could rely on distilled language features [49]. Further experiments, as well as discussions on limitations and future work, can be found on the project page.
@@ ACKNOWLEDGMENT
This work was funded by project SGR 00514 (Departament de Recerca i Universitats de la Generalitat de Catalunya) and CSIC project 202350E080 (ClothIRI). O.B. acknowledges travel support from ELISE (GA no 951847).
@@ A-B Language annotations
To enable consistent labeling across frames, we continuously track the NOCS coordinates of the garment mesh throughout the manipulation sequence. This ensures that the vertex assignments remain stable despite large deformations, which is crucial for accurate semantic labeling. An illustration of this tracking process is shown in Fig. 8, where a pair of pants is manipulated and the associated NOCS values are maintained across frames. Once the NOCS mapping is established, we apply axis-aligned thresholding to discretize the coordinate space and assign semantic labels to different garment regions, as shown in Fig. 9. This approach allows us to distinguish key parts such as sleeves, legs, and waistbands across diverse garment types using a unified representation. Notably, we do not threshold the front–rear direction, as it is not relevant for the folding actions considered in this work.
(a) Bring the right side of the trousers towards the left side and fold them in half.
(b) Bend the trousers in half, from bottom to top.
(c) Straighten out the bottom part of the trousers.
Figure 8: Continuous NOCS tracking across pick–and–place subactions. Each panel shows a single pick-and-place step rendered in two columns: the left image corresponds to the pre-action state, and the right to the post-action state. We overlay the user’s VR-tracked hand positions along with the corresponding text instruction. Unlike the vertex positions, the NOCS coordinates remain consistent throughout the manipulation, enabling stable vertex-to-semantic-region assignments as long as mesh topology is preserved.
@@
When annotating bimanual actions using NOCS coordinates, it is usual that the left and right pickers use different semantic locations, e.g., the left picker grabs the top right part, and the right picker grabs the bottom right part. In this case, we can infer that the common objective is to hold the right part of the garment, but one cannot trivially resolve many other situations. To do that, we designed a heuristics detailed in Algorithm 1, which outputs a common semantic location considering the positions of the left and right pickers given the context of the action.
@@ C-C Failures
Appendix D Limitations
D-A Previous approaches
@@
D-B BiFold dataset
@@ D-C BiFold model and evaluation
Among the baselines considered in this work, Foldsformer [3] and Deng et al. [5] use SoftGym as a simulator [36], while CLIPort [4] relies on PyBullet [54]. However, according to the benchmark performed by Blanco-Mulero et al. [46], the cloth physics of these simulators is inaccurate. Among all the simulators tested in the benchmark, MuJoCo [55] is chosen as the one achieving closer dynamics. Even if the pick and place positions are perfect, the success metric also considers errors from the manipulator. While simulators allow the grasp mesh vertices and hence achieve an ideal grasp, deploying on real robots introduces additional errors due to incorrect grasps or workspace limitations. This can be seen in some examples in simulation, and we include one of them in Fig. 28.
A natural next step could be to not rely on primitives and directly predict robot actions, which could help improve the grasp and evaluation of real robots. To do that we could use a diffusion policy [47] and condition it with the predicted pick and place actions [48].
In this work, we also focus on folding subactions instead of an end-to-end fold. However, performing a fold would amount to either following a predefined list of ordered instructions or using a LLM-based planner to break down the task into steps [45].